# GUARD: Guided Unlearning and Retention via Data Attribution for Large Language Models

### v1

@@ C.1 Key factors analysis for theoretical guarantee
    Unlearning rate $\eta$. An increased unlearning rate $\eta$ improves retention on the retain set ${\mathcal{D}}_{r}$ by amplifying gradient alignment (Lemma 4.1). However, it may slightly reduce unlearning effectiveness by overcorrecting the model on ${\mathcal{D}}_{f}$ (Lemma 4.2). Overall, a higher $\eta$ leads to a greater reduction in the Sacrifice Rate (Theorem 1).
@@
    Temperature $\tau$. Lowering the temperature induces a sharper attribution distribution across data points. This has three effects: (1) it enhances retention by emphasizing high-variance alignment directions (Lemma 4.1); (2) it may weaken unlearning effectiveness due to increased entanglement between ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$ (Lemma 4.2); and (3) it results in a larger reduction in the Sacrifice Rate, especially when $\delta_{\kappa}=\kappa/\tau\ll 1$ (Theorem 1).
@@
    Norm of the average forget gradient $\|\overline{g}_{f}\|$. Larger values of $\|\overline{g}_{f}\|$ have minimal effect on retention. However, they hinder unlearning by making the forget gradients more dominant, thereby resisting suppression (Lemma 4.2). This also leads to a smaller reduction in the Sacrifice Rate due to their influence on the denominator of the alignment term (Theorem 1).
Figure 2: Illustration of the standard unlearning pipeline for LLMs. The process consists of two main steps: (1) fine-tuning a pretrained LLM on a given dataset and (2) applying an unlearning algorithm to remove the influence of a specific subset of data.
Appendix D Algorithms
D.1 Update Rules of Baseline Methods
We list the baseline unlearning algorithms provided in TOFU benchmark [31], including Gradient Ascent (GA), Gradient Difference (GD), KL Minimization (KM), and Preference Optimization (PO).
@@ D.1 Update Rules of Baseline Methods
Gradient Difference (GD) [28]. The Gradient Difference method builds upon the gradient ascent approach by not only increasing loss on the forget set $D_{f}$ but also preserving model performance on the retain set $D_{r}$. The objective function for this approach is formulated as:
@@
Preference-Based Optimization. Inspired by Direct Preference Optimization (DPO) [37], this method substitutes responses in $D_{f}$ with alternative neutral responses such as “I do not have that information”. The loss function is as follows:
@@
D.2 GUARD objective for framework beyond Gradient Ascent
@@
D.3 GUARD algorithms for additional baselines
@@
2:Compute attribution scores $a_{i}$ for all $i\in I_{f}$triangleright Estimate data attribution for forget set
@@
5:     Compute unlearned model parameters $\theta_{GD}^{GUARD}$ using Eq. 26 $\triangleright$ GUARD Unlearning
@@
8:         Compute unlearned model parameters $\theta_{KM}^{GUARD}$ using Eq. 27 $\triangleright$ GUARD Unlearning
@@
11:              Compute unlearned model parameters $\theta_{PO}^{GUARD}$ using Eq. 30 $\triangleright$ GUARD Unlearning
12:         end if
13:     end if
@@
15:Output: Unlearned model weights $\theta_{B}^{GUARD}$.
Appendix E Additional results
E.1 Extra Settings
@@
All experiments were evaluated on equal-cost hardware, Ubuntu 22.04 LTS system with 2 Intel(R) Xeon(R) Gold 6138 CPUs with 20 cores, 4 NVIDIA Tesla V100-SXM2 GPUs with 32 GB memory, and 384 GB RAM. All experiments use 40 CPU cores and 4 GPUs. Unless otherwise specified, the following configurations are used: During the fine-tuning phase, we train for 5 epochs using a learning rate of $10^{-5}$ and a weight decay of 0.01. The batch size is set to 4 for Phi-1.5B and 1 for LLaMA-2-7B. During the data attribution phase, we use a threshold of $\tau=0.03$ unless stated otherwise. In the forgetting phase, by default, each experiment use 1 training epoch, a batch size of 1, a learning rate of $10^{-5}$, and a weight decay of 0.01.
E.2 Results for Phi-1.5B
We present results of Phi-1.5B as in Table 5, 6 and 7.
@@
Impact of Forgetting Proportion.
@@
E.3 Proxy for data attribution
@@
E.4 Hyperparameter Sensitivity
We present the unlearning effect of GUARD in Fig. 4, where we use the example of unlearning Phi-1.5B on TOFU (Forget Split = 1%). In Fig. 4, the x-axis use $-\log(\tau)$ to depict various $\tau$ for unification. Fig. 4 shows the performance across three evaluation metrics: ROUGE-1, ROUGE-1 drop $\Delta_{\text{ROUGE-1}}$, and Sacrifice Rate $\rho^{ul}$. The results demonstrate that our method significantly outperforms the baseline in maintaining the desired knowledge.
Figure 4: Hyperparameter Sensitivity Analysis

### v2

@@ D.1 Analysis of Key Factors for Theoretical Guarantees
    Unlearning rate $\eta$. An increased unlearning rate $\eta$ improves retention on the Retain Set ${\mathcal{D}}_{r}$ by amplifying gradient alignment (Lemma 1). However, it may slightly reduce unlearning effectiveness by overcorrecting the model on ${\mathcal{D}}_{f}$ (Lemma 2). Overall, a higher $\eta$ leads to a greater reduction in the Sacrifice Rate (Theorem 3).
@@
    Temperature $\tau$. Lowering the temperature induces a sharper attribution distribution across data points. This has three effects: (1) it can enhance retention by emphasizing high-variance alignment directions (Lemma 1); (2) it may weaken unlearning effectiveness due to increased entanglement between ${\mathcal{D}}_{f}$ and ${\mathcal{D}}_{r}$ (Lemma 2); and (3) it can result in a larger reduction in the Sacrifice Rate, especially when $\delta_{\kappa}=\kappa/\tau\ll 1$ (Theorem 3).
@@
    Norm of the average forget gradient $\|\overline{g}_{f}\|$. Larger values of $\|\overline{g}_{f}\|$ have minimal effect on retention. However, they hinder unlearning by making the forget gradients more dominant, thereby resisting suppression (Lemma 2). This also leads to a smaller reduction in the Sacrifice Rate due to their influence on the denominator of the alignment term (Theorem 3).
Figure 2: Illustration of the standard unlearning pipeline for LLMs, which is applying an unlearning algorithm to remove the influence of a specific subset of data (Forget Set).
Appendix E Algorithms
E.1 Update Rules of Baseline Methods
We list the baseline unlearning algorithms provided in TOFU benchmark Maini et al. (2024b), including Gradient Ascent (GA), Gradient Difference (GD), KL Minimization (KM), and Preference Optimization (PO).
@@ E.1 Update Rules of Baseline Methods
Gradient Difference (GD) Lu et al. (2022). The Gradient Difference method builds upon the gradient ascent approach by not only increasing loss on the forget set $D_{f}$ but also preserving model performance on the retain set $D_{r}$. The objective function for this approach is formulated as:
@@
Preference-Based Optimization. Inspired by Direct Preference Optimization (DPO) (Rafailov et al., 2023), this method substitutes responses in $D_{f}$ with alternative neutral responses such as “I do not have that information”. The loss function is as follows:
@@
E.2 GUARD objective for frameworks other than GA
@@
  • GUARD objective under Negative Preference Optimization.
    $L_{\text{NPO}}^{GUARD}=L^{GUARD}_{D_{f},\text{NPO}}(\theta):=\frac{1}{n_{f}}\sum_{i\in I_{f}}\omega^{GUARD}_{i}\cdot\frac{2}{\beta}\log\left(1+\left(\frac{\pi_{\theta}(y_{i}|x_{i})}{\pi_{\text{ref}}(y_{i}|x_{i})}\right)^{\beta}\right).$ (30)
E.3 GUARD algorithms for additional baselines
@@
2:Compute attribution scores $a_{i}$ for all $i\in I_{f}$ $\triangleright$ Estimate data attribution for forget set
@@
5:  Compute unlearned model parameters $\theta_{\text{GD}}^{\text{GUARD}}$ using Eq. 25 $\triangleright$ GUARD Unlearning
@@
8:   Compute unlearned model parameters $\theta_{\text{KM}}^{\text{GUARD}}$ using Eq. 26 $\triangleright$ GUARD Unlearning
@@
11:     Compute unlearned model parameters $\theta_{\text{PO}}^{\text{GUARD}}$ using Eq. 29 $\triangleright$ GUARD Unlearning
12:   else if $B$ is NPO then $\triangleright$ For Negative Preference Optimization
13:     Compute unlearned model parameters $\theta_{\text{NPO}}^{\text{GUARD}}$ using Eq. 30 $\triangleright$ GUARD Unlearning
@@
15:  end if
16:end if
17:Output: Unlearned model weights $\theta_{B}^{GUARD}$.
Appendix F Additional results
F.1 Additional Settings
@@
All experiments were evaluated on equal-cost hardware, Ubuntu 22.04 LTS system with 2 Intel(R) Xeon(R) Gold 6138 CPUs with 20 cores, 4 NVIDIA Tesla V100-SXM2 GPUs with 32 GB memory, and 384 GB RAM. All experiments use 40 CPU cores and 4 GPUs. Unless otherwise specified, the following configurations are used:
TOFU benchmark: During the fine-tuning phase, we train for 5 epochs using a learning rate of $10^{-5}$ and a weight decay of 0.01. The batch size is set to 4 for Phi-1.5B and 1 for LLaMA-2-7B. During the data attribution phase, we use a threshold of $\tau=0.03$ unless stated otherwise. In the forgetting phase, by default, each experiment use 1 training epoch, a batch size of 1, a learning rate of $10^{-5}$, and a weight decay of 0.01.
MUSE benchmark: MUSE provides pre-finetuned large models, which can be directly downloaded from its Huggingface repository. During the data attribution phase, we use a threshold of $\tau=0.03$ unless stated otherwise. In the forgetting phase, by default, each experiment use exactly the same setting compare to MUSE benchmark’s instruction.
F.2 Results for LLaMA-2-7B on the TOFU Dataset
We present the results in Table 45.
Table 4: Unlearning experiments with LLaMA-2-7B on 1% of all data points in TOFU. GUARD consistently reduces the Sacrifice Rates across all baselines and evaluation metrics, especially on the Retain Set. It can also lead to negative Sacrifice Rates (i.e., performance gains) on generalization sets, suggesting enhanced generalization via reduction of overfitting on the forget set.
$\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$
GA 87.87 91.25 289.06 58.15 73.91 172.38 31.54 64.55 119.48
GA + GUARD 49.82 55.32 124.73 39.59 52.36 90.89 17.68 -18.82 -5.93
KM 82.45 79.01 134.99 41.42 18.10 129.15 5.96 19.09 59.17
KM + GUARD 43.42 45.08 24.69 27.16 0.18 55.38 -2.94 -7.08 -1.13
PO 24.59 47.38 94.13 16.39 14.87 61.95 -15.34 5.84 -21.89
PO + GUARD 9.36 31.52 15.99 6.42 4.55 51.36 -17.40 -2.22 -29.21
GD 11.97 12.15 91.65 2.63 4.49 -15.86 -14.29 -17.52 -29.21
GD + GUARD 1.64 8.98 15.51 -1.53 -3.78 -18.04 -17.91 -18.59 -32.02
Table 5: Unlearning experiments with LLaMA-2-7B on 5% of all data points in TOFU. GUARD consistently reduces the Sacrifice Rates across all evaluation metrics and baselines, with particularly large improvements over baselines that exhibiting weak retention (the most substantial gains are observed against GA, which has the highest Sacrifice Rates).
$\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$
GA 90.57 118.59 401.01 80.77 88.02 313.25 64.76 37.84 216.97
GA + GUARD 64.31 87.93 237.29 61.73 65.49 236.95 36.65 5.32 165.84
KM 86.15 112.50 337.22 78.74 42.56 180.75 20.50 20.25 98.92
KM + GUARD 60.60 86.79 201.89 58.10 25.17 134.08 -5.24 -11.51 66.66
PO 47.96 104.14 200.59 42.45 25.22 152.90 6.24 -9.37 19.02
PO + GUARD 25.24 82.39 118.83 25.92 13.25 97.87 -4.73 -12.18 -11.29
GD 40.97 79.47 114.25 21.40 -14.44 -17.61 -21.13 -11.53 -11.50
GD + GUARD 18.71 59.30 86.16 10.49 -18.52 -23.77 -27.13 -13.95 -23.74
F.3 Results for Phi-1.5B on the TOFU Dataset
We present the results of Phi-1.5B in Tables 6, 7 and 8.
@@ F.2 Results for LLaMA-2-7B on the TOFU Dataset
Impact of Forgetting Percentage.
@@
F.4 Results for ICLM-7B on the MUSE BOOKS Dataset
Table 9: Unlearning experiments with ICLM-7B on the MUSE BOOKS dataset.
Methods Forget Set VerbMem Forget Set KnowMem PrivLeak Retain Set KnowMem
($\downarrow$) ($\downarrow$) ($\in[-5\%,5\%]$) ($\uparrow$)
BOOKS
w/o Unlearn 99.8 59.4 57.5-57.5 66.9
GA 0.0 0.0 $\mathbf{-25.0}$ 0.0
GA 0.0 0.0 31.3-31.3 10.3
KM 15.4 22.5 $\mathbf{-39.1}$ 35.9
KM + GUARD 16.0 20.3 42.2-42.2 47.2
GD 0.0 0.0 29.1-29.1 11.6
GD + GUARD 0.0 0.0 $\mathbf{-26.6}$ 28.1
NPO 0.0 0.0 -32.3 24.5
SimNPO 0.0 0.0 -22.6 47.8
FLAT (TV) 0.0 0.0 -45.2 28.9
SimImP 0.0 0.0 -26.4 22.5
NPO + GUARD 0.0 0.0 -38.16 52.71
We present the results of ICLM-7B on MUSE BOOKS dataset in Table 9.
F.5 Computational Efficiency Analysis
Table 10: Runtime of GUARD Unlearning for Single Epoch
Dataset Model Forget Proportion/Set Runtime of GA (min) Runtime of GA + GUARD (min)
TOFU LLaMA-2-7B 1% 4.2 5.1
5% 14.4 16.6
10% 32.3 35.5
Phi-1.5B 1% 2.8 3.0
5% 6.1 6.8
10% 14.5 15.6
MUSE LLaMA-2-7B NEWS 83.7 91.2
ICLM-7B BOOKS 116.0 127.9
To evaluate the computational overhead introduced by GUARD, we conducted runtime comparisons between standard Gradient Ascent (GA) and GA augmented with GUARD across different datasets and model architectures. Table 10 presents the runtime measurements for single epoch unlearning experiments conducted on our experimental setup with 4 NVIDIA Tesla V100-SXM2 GPUs.
The runtime comparison between GA and GA + GUARD provides a representative assessment of GUARD’s computational overhead. Importantly, the additional time introduced by GUARD primarily stems from the data attribution computation phase, which is independent of the choice of baseline unlearning method. This means that the overhead observed with GA would be similarly applicable when GUARD is integrated with other baseline methods.
A critical aspect of GUARD’s computational efficiency lies in the nature of its data attribution computation. The attribution score calculation described in Eq. 2 serves as a preprocessing step that is performed only once for each model parameters-dataset combination. Crucially, this computation is independent of the number of unlearning epochs and the frequency of unlearning operations. Once the attribution scores are computed for a given forget set, they can be reused across multiple unlearning sessions or when experimenting with different hyperparameters, amortizing the computational cost over multiple uses.
In conclusion, GUARD demonstrates practical computational efficiency with minimal time overhead. The one-time nature of the attribution computation, combined with the modest runtime increases observed across diverse experimental settings, establishes GUARD as a computationally viable enhancement to existing unlearning methods without imposing prohibitive computational costs.
F.6 Proxy for data attribution
@@
F.7 Hyperparameter Sensitivity
We present the unlearning effect of GUARD in Fig. 4, where we use the example of unlearning Phi-1.5B on TOFU (Forget Split = 1%). In Fig. 4, the x-axis use $\log(\tau)$ to depict various $\tau$ for unification. Fig. 4 shows the performance across three evaluation metrics: ROUGE-1, ROUGE-1 drop $\Delta_{\text{ROUGE-1}}$, and Sacrifice Rate $\rho^{ul}$. The results demonstrate that our method significantly outperforms the baseline in maintaining the desired knowledge.
Figure 4: Hyperparameter Sensitivity Analysis.